JEF meeting, December 20, 2024

From GlueXWiki
Jump to: navigation, search
Connection =

Meeting time and location:


Agenda

Preparation of parts for ECAL -- Sasha

-Four boxes for HV distribution boards have arrived to the lab. PCBs with HV connectors were installed on 2 distribution boards [1][2]
-Updated design of the LMS light source with 25 LEDS. PCB layout under preparation, expect to order the new PCB during next week.
-Four temperature sensors (RTD) arrived to the lab. They will be installed in the ECAL divider region during next week.
-Two additional fans for cooling the ECAL dividers arrived. Will be installed today.

-Installation of the FCAL2 detector & integration to GlueX -- Sasha, All

  • Implementing ECAL into the GlueX DAQ [3]
- rocecal3 crate was fixed [4] (initial DAQ implementation [5])
- implementing flash ADC scalers [6][7]
- translation table for the ECAL (finalize/check translation DB, generate xml file, import to the CCDB) - done [8]
- implementing and verifying ECAL LMS trigger [9]
- implementing and verifying the ECAL physics trigger. First event reconstructed using the GlueX software [10]
- Start implementing VTP trigger modules. Requires modification of firmware. The trigger crates may be unavailable for 1 - 1.5 months.
- The VTP was installed in the D2-1-TOP crate, IP address assigned, and can be booted now
- New version of DAQ, VTP included as a DAQ component [11]
  • Installation of scintillating paddles for cosmic tests [12] [13][14]
  • ECAL HV interlock check [15]
  • Cable map of the ECAL LG modules after installing ECAL [16]
  • Reinstalling cables on the ECAL LV distribution board [17]
  • Safety briefings - Vlad, Sasha
  • Detector installation page
  • Commissioning north part of the ECAL (instrumented with 2 temporary HV distribution boards) [18][19][20][21][22]
  • Commissioning of the south bottom region [23]
  • Commissioning of the south top region [24]
  • Commissioning of the top north region [25]
  • Commissioning of the bottom north region [26]
  • All HV distribution boards installed, set HV on the whole detector [27]
  • Fix LV issues (north top region) [28]
  • First cosmic event in the ECAL [29]
  • Temperature scan in the ECAL divider region [30][31][32]/4321188[33]
Meeting with Tim, additional fans will be installed
  • Repairing a few ECAL LMS optical fibers [34][35]
  • Preparing for the installation of the insulation on the detector face
- Removing Tedlar from the face of the detector [36]
- ECAL flash ADC scalers before removing detector face cover [37]
  • Preparing for the run -- Sasha
Page for shift Takers
ECAL in the Experimental Safety Assessment Document (ESAD)
EPAS for Maintenance of the Hall D ECAL (Jlab-PTW-5236)
- Installing additional fans (will be installed today)
- Installing insulation on the detector's face (done)

Restoring PS test setup [38]

ECAL commissioning -- Sasha, All

Laveen update[[39]]

Calibrations

-First look on cosmic data -- Zhikun
-Adjusting HVs

Analysis/Simulation

Minutes

Attendees: Bill, Phil, Igal, Liping, Sasha, Vlad, Zhikun, Jane, Olga, Laveen, Jane, Zisis

  • HV Distribution Boards:
    • installed 2 HV distribution boards on the north side
    • cables arranged and connected; issues resolved (by Sammy):
      • two connectors with broken lock clips
      • two connectors with small holes
      • multi-pin connectors being pressed by long screws
      • one faulty HV channel on the bottom north PCB—reconnected to a spare channel
  • Detector power-up and testing using the LMS:
    • entire detector powered up
    • fixed minor issues, such as loose PMT socket connections
    • confirmed all ECAL channels are operational
    • need to restore HV cable mapping in CSS after final board installation
  • Temperature sensors
    • two sensors on the top corners of the detector are operational (critical temperature region)
    • some issues with two other sensors, one positioned around the beam pipe, the other one on the bottom north corner (need to uninstall and fix)
    • two sensors will be included into the HV interlock (consider to include the fan into the interlock as well)
    • will be ready to collect cosmic ray data after that
  • Chris is finalizing the LMS light source design (to be ordered by end of 2024)
  • Bobby 3D printed parts for reference PMTs
  • ECAL calibration discussions to begin in the next JEF meeting.